OpenStudy (anonymous):

Solve -4(x + 10) - 6 = -3(x - 2). -40 -46 -52 52

jagr2713 (jagr2713):

Well first we distribute

jagr2713 (jagr2713):

so this becomes -4x-40-6=-3x+6

jagr2713 (jagr2713):

then we combine like terms -4x-46=-3x+6

jagr2713 (jagr2713):

Then we combine like terms again -4+3=-1 -x-46=6 add 46 on both sides -x=52 then we divide and get what ?

OpenStudy (anonymous):

Gahhh, I just totally got lost!

OpenStudy (anonymous):

I'm not sure what I'm dividing.

jagr2713 (jagr2713):

-x divided by 52 -x is another word for -1

OpenStudy (anonymous):

So it's still 52?

jagr2713 (jagr2713):

Caution! NEGATIVE divided by a Positive makes what?

OpenStudy (anonymous):

Oh wait.. -52?

OpenStudy (anonymous):

I really need help..

jagr2713 (jagr2713):

Correct!

OpenStudy (anonymous):

Can you help with another? :/

jagr2713 (jagr2713):

Sorry for the long wait. It didnt inform you you messaged me.

jagr2713 (jagr2713):

Yes

OpenStudy (anonymous):

Solve for L: A = L • W • H L = A times W over H

OpenStudy (anonymous):

and Below are the steps to solve an equation: Step 1: |x - 5| + 2 = 5 Step 2: |x - 5| = 5 - 2 Step 3: |x - 5| = 3 Which of the following is a correct next step to solve the equation? x + 5 = -3 -x - 5 = 3 x - 5 = -3 x + 5 = 3

OpenStudy (anonymous):

Oh wait, it's cut off. Solve for L: A = L • W • H L = A times W over H L = W times H over A L = A over quantity W times H L = W over quantity H times A

jagr2713 (jagr2713):

Wait i am confused which one is the question?

OpenStudy (anonymous):

I posted two questions, there first one is the LWAH one.

OpenStudy (anonymous):

do you know the answers?

jagr2713 (jagr2713):

This is how you'd solve the equation: Step 1: |x - 5| + 2 = 5 Step 2: |x - 5| = 5 - 2 Step 3: |x - 5| = 3 Step 4: x - 5 = 3 or x - 5 = -3

OpenStudy (anonymous):

and what about the other one?

jagr2713 (jagr2713):

L = A over quantity W times H